Write a system of equations that has (2, 7.5) as its solution.
Answer by venugopalramana(3286)

we need a system of eqns.,that has 2,7.5 as its solution.there are 2 solutions,hence we need to have 2 variables..let them be x and y
let us find their sum =x+y=2+7.5=9.5
let us find their difference=x-y=2-7.5=-5.5
hence the required system of eqns.are
x+y=9.5 and x-y = -5.5